# thanks: Contributed equallythanks: Contributed equally

### v3

@@ Appendix B Correlation functions and fluctuations
The pair distribution function $g(r,\phi,\theta)$ implies correlations and provides the probability density of finding a pair of particles at a distance $r$ with relative angles $\phi$ and $\theta$. The angle $\phi$ is between the relative position $r$ and the self-propulsion direction of the tagged particle, $\mathbf{\hat{e_{i}}}\cdot\mathbf{\hat{r_{ij}}}=\cos\phi$ and $\theta$ is the angle between the propulsion directions of the particle pair. To obtain this pair distribution function, histograms are created with the number of particles $N_{h}(r,\phi,\theta)$ at distance $r$, positional angle $\phi$ and relative angle $\theta$. The bins are chosen as $0.1b$ for $r$, $\pi/150$ for $\phi$ and $\pi/2$ for $\theta$. Then we need to normalize $N_{h}(r,\phi,\theta)$, to obtain the pair distribution function $g(r,\phi,\theta)$. The number of configurations is $t_{c}$($10000$ snapshots), over which histograms are made. The area of the annular segment of radial width $dr$, angular width $d\phi$ is given by $A_{r}=rdrd\phi$. The number density is $\rho=N/L^{2}$ and the size of the relative orientation bin is $d\theta$. Then the normalization factor is $N_{r}=\frac{A_{r}\rho Nt_{c}d\theta}{2\pi}$ and the pair distribution function is given by $g(r,\phi,\theta)=N_{h}(r,\phi,\theta)/N_{r}$.
@@
An example of this for isotropic phase is given in Fig. 6(a). The pair distribution functions are determined for the four different $\theta$ ranges as: $\theta$$0,\pi/2$], $\theta$$\pi/2,\pi$], $\theta$$\pi,-\pi/2$] and $\theta$$\pi/2,0$]. We see that in the isotropic phase, the probability of finding a neighboring particle is different in the front relative to the back of the direction of motion (orientation), indicative of the symmetry breaking in an active disordered gas [17, 38, 26]. In the main text, we show that $g(r,\phi)$ also distinguishes CLFs from CCFs, the former having a more quadrant-like structure compared to the latter’s hexatic structure. In addition, we compare the structure of $g(r,\phi)$ for the band phases in Fig. (6)(b), for the CLF and CCF density bands. Note that the band structure is more dense in the case of $\chi_{t}=0$, resulting a more intense correlation in $g(r,\phi)$. A related quantity that we compute is the radial $g(r)$, given by
$g(r)=\frac{1}{N}\sum_{i,j}\langle\delta\Big{(}|\mathbf{r}_{i}-\mathbf{r}_{j}|-r\Big{)}\rangle$ (23)
this is plotted in Fig. (3), distinguishing the CLF and CCF.
An additional feature of the CLF is that they have a steady-state persistent fluctuations of both density and polarization fields (dynamical steady-state - note that our simulations are effectively noiseless), which is absent in the CCF. In Fig.7, we show fluctuations for positions and orientations for CLF and CCF . Here, the Density fluctuations is given by: $\langle(\Delta\rho)^{2}\rangle=\langle\rho^{2}\rangle-\langle\rho\rangle^{2}$. Here, $\langle\quad\rangle$ denotes a spatial average. Note that the $\rho$ field has to be obtained via binning. We also measure Orientational (angular) fluctuations, which is given by $\langle(\Delta\theta)^{2}\rangle=\langle\theta^{2}\rangle-\langle\theta\rangle^{2}$. Here, $\langle\quad\rangle$ denotes average over colloids.
Figure 7: Comparison of CLF ($\chi_{t}=0$) and CCF ($\chi_{t}>0$) via density fluctuations ($\langle(\Delta\rho)^{2}\rangle$), and orientational fluctuations ($\langle(\Delta\theta)^{2}\rangle$). For the CLF, the panel on the left shows residual fluctuations, whereas for the CCF, the fluctuations die out in the steady-state. In the panel on the right, a similar phenomenon is seen for the orientational fluctuations. Note that both curves are strongly correlated (shared monotonicity).
Appendix C Description of the supplementary movies

### v4

@@ Appendix A Pair correlation functions
The pair distribution function $g(r,\varphi,\theta)$ implies correlations and provides the probability density of finding a pair of particles at a distance $r$ with self-propulsion angle $\theta$. Here, $\varphi$ is the angle between the relative position $r$ and the self-propulsion direction of the tagged particle, $\hat{\mathbf{e}}_{i}\cdot\hat{\mathbf{{r}}}_{ij}=\cos\varphi$ and $\theta$ is the angle between the propulsion directions of the particle pair. To obtain this pair distribution function, histograms are created with the number of particles $N_{h}(r,\varphi,\theta)$ at distance $r$, positional angle $\varphi$ and angle $\theta$. The bins are chosen as $0.1b$ for $r$, $\pi/150$ for $\varphi$ and $\pi/2$ for $\theta$. Then we need to normalize $N_{h}(r,\varphi,\theta)$, to obtain the pair distribution function $g(r,\varphi,\theta)$. The number of configurations is $t_{c}$($10000$ snapshots), over which histograms are made. The area of the annular segment of radial width $dr$, angular width $d\varphi$ is given by $A_{r}=rdrd\varphi$. The number density is $\rho=N/L^{2}$ and the size of the relative orientation bin is $d\theta$. Then the normalization factor is $N_{r}=\frac{A_{r}\rho Nt_{c}d\theta}{2\pi}$ and the pair distribution function is given by $g(r,\varphi,\theta)=N_{h}(r,\varphi,\theta)/N_{r}$.
@@
An example of this for isotropic phase is given in Fig. 8(a). The pair distribution functions are determined for the four different $\theta$ ranges as: $\theta$$0,\pi/2$], $\theta$$\pi/2,\pi$], $\theta$$\pi,-\pi/2$] and $\theta$$\pi/2,0$]. We see that in the isotropic phase, the probability of finding a neighboring particle is different in the front relative to the back of the direction of motion (orientation), indicative of the symmetry breaking in an active disordered gas [17, 46, 26]. In the main text, we show that $g(r,\varphi)$ also distinguishes CLFs from CCFs, the former having a four-lobe structure compared to the latter’s hexatic structure. In addition, we compare the structure of $g(r,\varphi)$ for the band phases in Fig. (8)(b), for the CLF and CCF density bands. We see that the CLF and CCF bands are distinguished strongly via this measure: in the former there is a strong likelihood of finding neighbouring colloids in either the front or back of the propulsion direction, whilst in the latter the aggregate structure resembles that of a (non-polar) liquid due to the presence of cross bands. Note that the band structure is more dense in the case of $\chi_{t}=0$, resulting a more intense correlation in $g(r,\varphi)$. A related quantity that we compute is the radial $g(r)$. This is plotted in Fig. (4), distinguishing the CLF and CCF.
Figure 8: Pair correlation function $g(r,\varphi,\theta)$ for the isotropic phase for four different $\theta$ ranges in panel (a). In panel (b), we show pair correlation function $g(r,\varphi)$ in the band phase for $\chi_{t}=0$ and $\chi_{t}>0$.
Appendix B Description of the supplementary movies
@@
  • Movie III. In the case of $\chi_{t}=0$ (CLF), particles form several density bands through out the space. Additionally, less dense waves (which move perpendicular to the direction of density bands) are also seen to form with these density bands. The fixed parameter values are: $L=418$, $b=1$, $dt=0.01$, $D_{r}=10^{-3}$, $D_{t}=10^{-3}$, $\kappa=175$, $v_{0}=50$, $\chi_{t}=0$, $\chi_{r}=2.0$.